A factory worker pushes a 30.1 kg crate a distance of 4.1 m along a level floor at constant velocity by pushing downward at an angle of 28 ∘ below the horizontal. The coefficient of kinetic friction between the crate and floor is 0.24.

1-What magnitude of force must the worker apply to move the crate at constant velocity?

2-How much work is done on the crate by this force when the crate is pushed a distance of 4.1m?

3-How much work is done on the crate by friction during this displacement?

4-How much work is done by the normal force?

5-How much work is done by gravity?

6-What is the total work done on the crate?
Answer #1

1] Let applied force be F. For constant velocity net horizontal force = 0

F*cos theta - friction = 0

F*cos 28 degree - u*(mg+F*sin 28 degree) = 0

   F*cos 28 degree - 0.24*(30.1*9.8+F*sin 28 degree) = 0

       F*cos 28 degree - 0.24*(30.1*9.8+F*sin 28 degree) = 0

0.882947*F -0.24*(30.1*9.8+F*0.46947) = 0

F = 91.9 N answer

2] Work done by applied force= F*s*cos theta

= 91.9*4.1*cos 28 degree

= 333 J answer

3] Work done by friction = - 333 J because friction is equal and opposite to horizontal force

4] work done by the normal force = 0 J as there is no displacement in vertical.

5] work done by gravity = 0J

6] Total work done = 0 J